A chemist has three acid solutions. The first contains 20% acid, the second contains 30% acid and the third contains 60% acid. She uses all three solutions to obtain a final mixture of 80 litres containing 40% acid. She uses twice as much of the 60% acid solution as the 20% solution. She wants to find out the number of litres of each solution she would use to obtain the final mixture.

Answer by lwsshak3(11628)

A chemist has three acid solutions. The first contains 20% acid, the second contains 30% acid and the third contains 60% acid. She uses all three solutions to obtain a final mixture of 80 litres containing 40% acid. She uses twice as much of the 60% acid solution as the 20% solution. She wants to find out the number of litres of each solution she would use to obtain the final mixture.
**
let x=litres of 20% solution to be used
2x=litres of 60% solution to be used
(80-x-2x)=(80-3x) litres of 30% solution to be used
..
20%x+60%*2x+30%(80-3x)=40%*80
.2x+1.2x+24-.9x=32
.5x=32-24=8
x=16
2x=32
80-3x=32
..
litres of 20% solution to be used=16
litres of 60% solution to be used=32
litres of 30% solution to be used=32